\begin{thebibliography}{36}
	\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi
	\expandafter\ifx\csname bibnamefont\endcsname\relax
	\def\bibnamefont#1{#1}\fi
	\expandafter\ifx\csname bibfnamefont\endcsname\relax
	\def\bibfnamefont#1{#1}\fi
	\expandafter\ifx\csname citenamefont\endcsname\relax
	\def\citenamefont#1{#1}\fi
	\expandafter\ifx\csname url\endcsname\relax
	\def\url#1{\texttt{#1}}\fi
	\expandafter\ifx\csname urlprefix\endcsname\relax\def\urlprefix{URL }\fi
	\providecommand{\bibinfo}[2]{#2}
	\providecommand{\eprint}[2][]{\url{#2}}
\bibitem{sm} See Supplementary Material for details.  
	
\bibitem{evoting}
L.~Fouard, M.~Duclos, and P.~Lafourcade, ``Survey on electronic voting
schemes.''

\bibitem{dsignature}
K.~Nguyen, F.~Bao, Y.~Mu, and V.~Varadharajan, ``Zero-knowledge proofs of
possession of digital signatures and its applications,'' in {\em Information
	and Communication Security} (V.~Varadharajan and Y.~Mu, eds.), vol.~1726 of
{\em Lecture Notes in Computer Science}, pp.~103--118, Springer Berlin
Heidelberg, 1999.

\bibitem{Fortnow}
L.~Fortnow, ``The complexity of perfect zero-knowledge,'' in {\em Proceedings
	of the Nineteenth Annual ACM Symposium on Theory of Computing}, STOC '87,
(New York, NY, USA), pp.~204--209, ACM, 1987.

\bibitem{Feige}
U.~Feige, A.~Fiat, and A.~Shamir, ``Zero-knowledge proofs of identity,'' {\em
	Journal of Cryptology}, vol.~1, no.~2, pp.~77--94, 1988.

\bibitem{HorodeckiZKP}
P.~{Horodecki}, M.~{Horodecki}, and R.~{Horodecki}, ``{Zero knowledge
	convincing protocol on quantum bit is impossible},'' {\em eprint
	arXiv:quant-ph/0010048}, Oct. 2000.

\bibitem{kentrel}
A.~{Kent} {\em Phys. Rev. Lett.}, 1999.

\bibitem{kentrelfinite}
A.~Kent, Secure Classical Bit Commitment using Fixed Capacity
Communication Channels, 
J. Cryptology  {\bf 18} (2005) 313-335.  

\bibitem{bcsummoning}
A.~{Kent}, ``{Unconditionally secure bit commitment with flying qudits},'' {\em
	New Journal of Physics}, vol.~13, no.~11, p.~113015, 2011.

\bibitem{AdlamKent1}
E.~{Adlam} and A.~{Kent}, ``{Deterministic relativistic quantum bit
	commitment},'' {\em International Journal of Quantum Information}, vol.~13,
p.~1550029, June 2015.

\bibitem{AdlamKent2}
E.~{Adlam} and A.~{Kent}, ``{Device-independent relativistic quantum bit
	commitment},'' {\em Physical Review A}, vol.~92, p.~022315, Aug. 2015.

\bibitem{Barnett}
S.~M. {Barnett}, A.~{Chefles}, and I.~{Jex}, ``{Comparison of two unknown pure
	quantum states},'' {\em Physics Letters A}, vol.~307, pp.~189--195, Feb.
2003.

\bibitem{ensemblecomparison}
M.~{Sedl{\'a}k}, M.~{Ziman}, V.~{Bu{\v z}ek}, and M.~{Hillery}, ``{Unambiguous
	comparison of ensembles of quantum states},'' {\em Physical Review A},
vol.~77, p.~042304, Apr. 2008.

\end{thebibliography}

\documentclass[aps,prl,twocolumn,groupedaddress]{revtex4}
\usepackage{amsfonts,amssymb}
\usepackage{amsthm}
\usepackage{tikz}
\usetikzlibrary{calc}
\usetikzlibrary{arrows}
\usepackage{tikz-3dplot}
\usepackage{color} 
\usepackage[fleqn]{amsmath}
\usepackage{amsthm}
\usepackage{amssymb}
\usepackage{amsfonts}
\usepackage{bbm}
\usepackage{epsfig}
\usepackage{times}
\usepackage{hyperref}
\usepackage{bm}
\usepackage{float} 
\usepackage{appendix} 
\usepackage{mathrsfs}
\usepackage{appendix}

\newcommand*{\braopket}[3]{\langle #1| #2 | #3\rangle}

\usepackage{graphicx}

\DeclareGraphicsExtensions{.png}

\newcommand{\comment}[1]{}
\newcommand{\ket}[1]{\left |  #1 \right\rangle}
\newcommand{\bra}[1]{\left \langle #1  \right |}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}

\theoremstyle{definition}

\begin{document}

	\title{Knowledge-Concealing Evidencing of Knowledge about a Quantum State}

\author{Emily \surname{Adlam}} \affiliation{Centre for Quantum
	Information and Foundations, DAMTP, Centre for Mathematical
	Sciences, University of Cambridge, Wilberforce Road, Cambridge, CB3
	0WA, U.K.}  \author{Adrian \surname{Kent}} \affiliation{Centre for
	Quantum Information and Foundations, DAMTP, Centre for Mathematical
	Sciences, University of Cambridge, Wilberforce Road, Cambridge, CB3
	0WA, U.K.}  \affiliation{Perimeter Institute for Theoretical
	Physics, 31 Caroline Street North, Waterloo, ON N2L 2Y5, Canada.}
	\date{\today}

\begin{abstract} Bob has a black box that emits a single pure
state qudit which is, from his perspective, 
uniformly distributed.   Alice 
wishes to give Bob evidence that she has knowledge about the emitted state
while giving him little or no information about it.
We show that zero-knowledge evidencing of such
knowledge is impossible in quantum relativistic protocols, 
extending a previous result of Horodecki et
al..  We also show that no such protocol can be both sound and complete. 
We present a new quantum relativistic protocol which we conjecture to be close
to optimal in security against Alice and which reveals little
knowledge to Bob, for large dimension $d$.  We
analyse its security against general attacks by Bob
and restricted attacks by Alice.   
\end{abstract}
	
		\maketitle
	
	\section{Introduction}
	
Zero-knowledge proving is a cryptographic primitive in which
one agent proves a fact to another agent without giving away
any information other than that the fact is true. It has a
wide range of practical applications, particularly in
electronic voting schemes \cite{evoting} and digital signature
schemes \cite{dsignature}, and is also used for a variety of
theoretical purposes, such as showing that a language is easy
to prove \cite{Fortnow}. Zero-knowledge proving of
\emph{knowledge}, where Alice is required to prove only that she
knows some fact, without giving Bob any information about the
fact itself, is a particularly useful version of this task
which plays a key role in a number of identification protocols
\cite{Feige}.
		
Horodecki et al.\@ \cite{HorodeckiZKP}
explored the possibility of what they called a ``zero knowledge convincing protocol
on quantum bit''. 
In their model, a verifier (henceforth called Bob) knows he has 
a single copy of a pure 
qubit, but has no other information about the state.   
A prover (henceforth called Alice)
wishes to make a prediction that Bob can verify and
that will hold with certainty
only if she knows what the state is,  
without giving Bob any additional information about its identity. 
They showed that no non-relativistic protocol involving classical
information exchanges and quantum Alice-to-Bob communications can 
implement this task securely \cite{HorodeckiZKP}.
They also discussed  
some protocols that implement very weak versions of the
task, either giving Bob a great deal of information about
the qubit, or giving him only weak evidence of Alice's knowledge,
or both.   

Horodecki et al.\@'s pioneering discussion was 
informal on some points. 
It did not fully distinguish cases in which Alice has classical
knowledge about Bob's quantum state (e.g. a classical data string
describing it) from cases in which she has quantum knowledge (e.g. a box able to make only some
fixed number of copies).     Nor did it underline that Alice cannot
{\it prove} that she knows a precise 
classical description of a single quantum state even if she is not
concerned about 
about giving Bob information.  This is because the classical information
about the state that can be extracted by measurement is bounded,
and Alice always has a boundedly nonzero chance of guessing this
information even if she knows nothing about the state.   (See Theorem
2 below.)  Alice may also have a high chance of guessing the information even
if she has only partial information about the quantum state -- for example, she can predict the outcome of a complete projective measurement
on a qubit with probability $\frac{1}{2}$. 
Similarly, even if she only knows a dimension $2$ subspace in which a qudit
lies, she can still specify a complete projective measurement whose outcome
she can predict with probability $\frac{1}{2}$. 
Moreover, if the state is $\eta$ and Alice believes it is $\eta'$,
where the fidelity $F(\eta, \eta')$ is close to $1$, then she is 
almost as likely to pass any protocol testing her knowledge of $\eta$ 
as she would be if she knew $\eta$.   
Since Bob only has a single copy of the state, Alice cannot provide
more evidence by repeating a protocol that tests her knowledge. 
By contrast, in classical contexts Alice's chances of success can typically be made 
arbitrarily small by iteration, so that one can
reasonably (modulo epsilonics) speak of classical zero knowledge
\emph{proofs}.  

Other interesting questions left open include: What can relativistic 
protocols achieve?  How much evidence can Alice provide?   
What bounds exist on the tradeoffs 
between the evidence Alice provides and the amount of
knowledge she gives away?  Are there protocols strong enough for
practical cryptographic purposes?   How do the answers depend
on the dimension $d$ of the state space?   

We explore these questions below, beginning 
with some formal definitions.  
We prove a stronger no-go theorem showing that 
no protocol that provides non-trivial evidence of Alice's 
knowledge about a pure quantum state of finite
dimension can prevent Bob from acquiring some additional
knowledge about the state, even in the setting of relativistic
quantum cryptography. 
We also prove a bound on the strength of evidence Alice can provide.   
Since proofs of knowledge of a finite-dimensional quantum state are
not possible, and zero-knowledge protocols that give some evidence
of knowledge are also not possible, we then consider the weaker
but feasible task of \emph{knowledge-concealing evidencing of knowledge about a quantum state}
(KCEKQS).  

A KCEKQS protocol requires 
Alice to give Bob evidence that she has some 
form of knowledge about a quantum state whilst giving him 
incomplete information about the state.  
Ideally, a successfully completed protocol should
give Bob as much evidence as possible, without assuming Alice's honesty.
Ideally, too, the protocol should ensure as small a bound as possible on the
information obtainable by Bob, whether or not it is successfully
completed or he honestly follows it.    We discuss some simple protocols, generalising 
protocols previously considered by Horodecki et
al. \cite{HorodeckiZKP}, and show that they are relatively weak 
in knowledge-concealment, or in evidencing knowledge, or both.  
We then propose a new relativistic quantum protocol. 
We show it is secure against restricted attacks by Alice and
general attacks by Bob, for large $d$, in a sense we make precise below.
We conjecture this remains true for general attacks by
Alice.
	
\section{Definitions}

We assume for now that Alice has no option to 
abort the protocol; allowing an abort option does not significantly 
change our main results \cite{sm}. 
	
A non-relativistic \emph{knowledge-concealing evidencing of knowledge about a quantum state
(KCEKQS)} protocol involves two mistrustful parties, Alice and Bob,
occupying disjoint secure laboratories.  We assume that each party has trusted error-free
devices in their own laboratory; both parties trust error-free classical and
quantum communication channels between the laboratories; we consider 
errors and losses later \cite{sm}.  
Bob begins in possession of a quantum system $Q_B$
known to be prepared in some pure state $\eta = \ket{\eta} \bra{\eta}$ drawn uniformly 
at random from $Q_B$.   The protocol 
requires Alice and Bob to act on alternate rounds and terminates
after a fixed finite number of rounds. 
Each round may require a party to carry out unitary operations
and/or measurements on a quantum system 
in their possession and/or to send classical and/or quantum
information to the other party.   The protocol may specify that these
actions are probabilistically determined, according to given
probability distributions.     
The final round of the protocol requires Bob to generate one of two possible outcomes,
$0$ and $1$, from the classical and quantum information in his
possession.   These correspond to Bob's rejecting or accepting that 
Alice has provided evidence of knowledge about $\eta$.    
We write $p(0)$ and $p(1)$ for the outcome probabilities.  

In a relativistic KCEKQS protocol, each party may have several trusted
agents occupying separate secure laboratories, with secure
communications between them, lying within pre-agreed regions. 
One agent of Bob's initially possesses $Q_B$.   The protocol
requires Alice's and Bob's agents to carry out unitary operations
and/or measurements on quantum systems in their possession and to send classical and/or quantum
communications to given other agents of the same party and/or the other
party, within their agreed location regions and within agreed time intervals.   
The protocol may specify these
actions are probabilistically determined, according to given
probability distributions.     
The protocol terminates after a fixed finite number of such actions.
The final prescribed action is for one of Bob's agents to 
generate one of two possible outcomes,
$0$ and $1$, from the classical and quantum information in his
possession, as above. 

We will characterise the efficiency of an ideal KCEKQS protocol by three parameters
$\epsilon_{\rm C}$, $\epsilon_{\rm K}$ and $\epsilon_{\rm S}$; we 
discuss other relevant features of KCEKQS protocols in the
supplementary material \cite{sm}.   
When evaluating these parameters for specific protocols, we will
mostly consider the ideal case of error-free devices and channels.  
In realistic implementations, channel noise, device
errors and losses may alter the parameter values.\footnote{Typically, 
uncorrected noise, errors and losses will increase $\epsilon_{\rm C}$ and decrease
$\epsilon_{\rm K}$, and uncorrected losses will decrease $\epsilon_{\rm S}$.}   
However, our no-go theorems below still hold in reasonable models of
noise, errors and losses, so long as these are uncorrelated with
$\eta$ and with any knowledge Alice may have of or about $\eta$.  

We define our parameters by the following criteria, in each case
averaging
over $\eta$: 
\begin{itemize}
			
\item \textbf{Completeness:} If Alice has a precise classical
description of $\eta$ and both parties perform the protocol correctly, then $p(1) = 
1 - \epsilon_{\rm C}$.   
			
\item \textbf{Soundness:} If Alice has no classical or quantum
information about the state $\eta$, then $\epsilon_{\rm S}$ is 
the supremum of $p(1)$ over all possible (honest or dishonest)
strategies for Alice, assuming that Bob performs the protocol correctly.  
			
\item \textbf{Knowledge-concealing:} Suppose that Alice performs the protocol
correctly and at the start of the protocol Bob knows
nothing about the state $\eta$.    Then $\epsilon_{\rm K}$ is the supremum
of the expected squared fidelity $F^2 ( \eta, \phi ) = | \, \langle \eta | \phi
\rangle \, |^2$ over (honest or dishonest) strategies that give Bob the value
of a pure state $\phi$ as a guess for $\eta$.  

\end{itemize} 			

Let $\epsilon_{\rm M}$ be the
supremum of the same expected squared fidelity obtainable by Bob if he does not 
take part in the protocol and carries out quantum operations and
measurements on $Q_B$.  
We call $\epsilon_{\rm K} - \epsilon_{\rm M}$ the \emph{knowledge
  gain} available from the protocol to a dishonest Bob.  
We say the protocol is \emph{ zero-knowledge}
if $\epsilon_{\rm K} = \epsilon_{\rm M}$,   We say it is \emph{non-trivial} if $1 - \epsilon_{\rm C} > \epsilon_{\rm S}$.   
			
As defined above, general protocols allow both classical and quantum
communications in both directions.   We will also consider 
examples with more restricted communications. 
Extending the discussion of Ref. \cite{HorodeckiZKP}, we consider
\emph{classical} protocols, in which Alice and Bob employ only
classical communication, \emph{quantum A-to-B}
protocols, which additionally allow quantum communications from A to
B, and similarly \emph{quantum B-to-A} protocols. 
	
\section{No-go theorems}

Horodecki et al.\@$~$\cite{HorodeckiZKP} showed that no non-relativistic 
KCEKQS classical or quantum A-to-B protocol for an unknown qubit 
has $\epsilon_{\rm C} =0$, $\epsilon_{\rm
  S}<1$, and is zero-knowledge.    
We establish here a considerably more general result, applying to 
relativistic KCEKQS protocols for qudits with two-way classical
and/or quantum communications, 
and general parameter values.   Our statements apply
to protocols whose security is based only on quantum theory and
special relativity, i.e. within the standard scenario for
unconditionally secure relativistic quantum cryptography \cite{kentrelfinite}. 

\begin{theorem} 

There exists no non-trivial zero-knowledge KCEKQS protocol. \cite{sm}

\end{theorem} 

We also establish a tradeoff between completeness and soundness which 
bounds the degree of evidence Alice can provide:
 
\begin{theorem} 
For any qudit KCEKQS protocol, $\frac{\epsilon_{\rm S}}{1 -
  \epsilon_{\rm C}} \geq \frac{1}{d}$.  \cite{sm}
\label{zeroknowledgetheorem}
\end{theorem}

In particular, theorem \ref{zeroknowledgetheorem} means that for small $d$, $\epsilon_{\rm S}$ and $\epsilon_{\rm C}$ cannot both be close
to $0$, regardless of the value of $\epsilon_{\rm K}$.
This makes the case of large $d$ particularly interesting to explore.

We observe that the bound of theorem \ref{zeroknowledgetheorem}
is tight. 
For 
example, it is attained by a protocol in which Alice predicts to Bob
the outcome of a projective measurement that includes $\eta$ on
the system $Q_B$: this has $\epsilon_{\rm S} = \frac{1}{d}$ and
$\epsilon_{\rm C}=0$.  More generally, it is attained for a protocol
in which Alice is required to predict this outcome and also predict the
outcome of some independent random event with success probability $p$:
this has $\epsilon_{\rm S} = \frac{p}{d}$ and $\epsilon_{\rm C}=
1-p$. 
We say a KCEKQS qudit protocol is {\it CS-optimal} if $\epsilon_{\rm S} = \frac{1}{d},
\epsilon_{\rm C} = 0$. 

\section{Protocols }
	
\subsection{Classical A-to-B \label{classicalmessage}}

As noted above, Horodecki et al.\@ \cite{HorodeckiZKP} argue that
non-trivial non-relativistic zero-knowledge classical A-to-B 
protocols with $\epsilon_{\rm C} = 0$ are impossible for a qubit.
In any such protocol, Alice must predict some measurement outcome,
and any measurement prediction that holds with certainty for 
a pure qubit $\eta$ and is not certain for a random qubit 
allows Bob to identify $\eta$ exactly,
and so has $\epsilon_{\rm K} = 1$. 

One might instead consider protocols with
$\epsilon_{\rm C} > 0$, in which Alice chooses a projective 
measurement which includes a randomly chosen projector $P$ from those
with $\braopket{\eta}{P}{\eta} = 1 - \epsilon_{\rm C}$.
One might also consider strengthening such protocols
by allowing Alice to use a secure relativistic bit commitment
\cite{kentrel,kentrelfinite, bcsummoning, AdlamKent1, AdlamKent2}
to commit her predicted outcome, unveiling this commitment if and
only if Bob's reported outcome agrees with her prediction.
However, such protocols still have either $\epsilon_{\rm K}$ or 
$\epsilon_{\rm C}$ large \cite{sm}. 
			
\subsection{Quantum A-to-B \label{quantummessage}}

Horodecki et al.\@ \cite{HorodeckiZKP} also consider a protocol where
Alice gives Bob a copy of $\eta$; as they note, such protocols
can achieve 
$\epsilon_{\rm C} = 0$ and 
$\epsilon_{\rm K} < 1$.
Indeed it is possible to achieve $\epsilon_{\rm K} \ll 1$ for large $d$.  In this
sense, for large $d$, their protocol outperforms the classical A-to-B protocols
just discussed. 
However, one needs to consider the tradeoffs
between $\epsilon_{\rm K}, \epsilon_{\rm C}$ and $\epsilon_{\rm S}$.  
It is also worth highlighting that this protocol requires
Alice only to possess quantum information about $\eta$ rather
than classical information.   Alice can ensure $p(1)=1$  
even if she only has a black box that will make only a single copy 
of $\eta$ and has no other classical or quantum information about
$\eta$.    
	
We extend the discussion of Ref. \cite{HorodeckiZKP}
by considering a generalisation of their protocol in which
Alice gives Bob $N$ copes of $\eta$: 

\begin{enumerate} 
	
	\item 	Alice prepares $N$ systems $\{ S_i \}$ in the state $\eta$ and gives them to Bob. 
	
	\item Bob performs a measurement $\{ \Pi_S, \mathbb{I} - \Pi_S\}$
where $\Pi_S$ is the projector onto the symmetric
subspace of the joint state space of the $\{ S_i \}$ and
$Q_B$.\footnote{A motivation for this choice is that, given`a system in
state $\psi^{\otimes n}$ and another system in state
$\phi^{\otimes m}$, for some integers $m, n$, the measurement
$\{ \Pi_S, \mathbb{I} - \Pi_S\}$ (i) always gives outcome $1$ if
$\psi = \phi$ (ii) maximises the probability of outcome $0$ if $\psi
\neq \phi$, among measurements satisfying (i) \cite{Barnett, ensemblecomparison}.}
	
	\item If the result is $\Pi_S$, Bob accepts; otherwise he
          rejects. 
 
\end{enumerate} 

If Alice knows $\eta$ and follows the protocol, Bob will accept, so
this protocol achieves $\epsilon_{\rm C} = 0$. 
We show \cite{sm} that
 $\epsilon_{\rm S}=\frac{1}{N + 1} + \frac{N}{d(N + 1)}$.
So, for $N=1$, we have $\epsilon_{\rm S} = \frac{1}{2} +
\frac{1}{2d}$, while, for $N$ large,  $\epsilon_{\rm S} \rightarrow \frac{1}{d}$, 
so the protocol tends to CS-optimality in this limit.
However, we also show \cite{sm} that $\epsilon_{\rm K} = 
\frac{N + 2}{N + d+1}$, which tends to $1$ for large $N$,
while $\epsilon_{\rm M} = \frac{2}{d+1}$.   
Thus $\epsilon_{\rm K} >  \frac{1}{d
\epsilon_{\rm S}}$  and $\epsilon_{\rm K} - \epsilon_{\rm M} >
\frac{d-1}{d+1} \frac{N}{N+2} \frac{1}{d \epsilon_{\rm S}}$ 
for all $d,N$, which are relatively poor tradeoffs. 
In particular, near 
CS-optimality ($ \epsilon_{\rm S} \approx \frac{1}{d}$) 
implies near-zero concealment ($\epsilon_{\rm K} \approx 1$)
and implies significant knowledge gain ($\epsilon_{\rm K} - \epsilon_{\rm M}
\approx \frac{d-1}{d+1}$).   
 
\subsection{A quantum B-to-A protocol \label{quantmvp}} 

We now propose a new relativistic protocol that involves quantum B-to-A
and two way classical communication.   
Given certain security assumptions, 
we show that,
the protocol asymptotically tends to CS-optimality in the 
limit where the security parameter $N$ is large, taking
the parameter $q = \lceil \frac{N}{d} \rceil $.
We conjecture that this remains true without the relevant assumptions.

We show that $ \epsilon_{\rm K} \leq \frac{4}{d+1}$, while 
$\epsilon_{\rm M} = \frac{2}{d+1}$, so 
$\epsilon_{\rm K} \rightarrow 0$ and 
$\epsilon_{\rm K} - \epsilon_{\rm M} \rightarrow 0$ 
for large $d$ \cite{sm}. 

\begin{enumerate} 
	
\item Alice and Bob agree in advance on positive integer
security parameters $N$ and $q$. 
	
\item Bob prepares $N $ quantum systems $\{ S_i \}$ in states chosen
uniformly at random.
	
\item Bob randomly permutes the systems $\{ S_i \}$ and the system
$Q_B$, assigns them all indices from $1$ to $N + 1$, and then gives
all $N + 1$ systems, labelled by their indices, to Alice.
	
\item Alice carries out the projective measurement
    $\{ \eta  \, , \,  \mathbb{I} -  \eta \}$ on each of the $N + 1$ systems that
  Bob gave her.   Write $C'$ for the list of indices for
which she obtains outcome $ \eta$; 
let $| C' | = q'$.  If $q' \leq q$, she forms a list $C = C' \cup D$, 
where $D$ is a list of $(q-q')$ copies of the dummy index $0$.\footnote{This dummy index
prevents cheating strategies in which Bob uses the number of Alice's
commitments, made at the next step, to extract additional 
information about the state.}.  If $q' > q$, she picks a random size $q$
sublist $C$ of $C'$.   
	
\item Alice randomly permutes $C$ and then
performs $q$ relativistic bit string commitments \cite{sm} committing her
to each of the indices in the permuted list.  Each bit string commitment is set up
so that Alice can commit to any index in $\{0, 1, 2,  \ldots N+1\}$.   
	
\item Bob tells Alice the index $x \in \{ 1, \ldots, N+1 \}$ that he assigned to $Q_B$.
	
\item If $x \in C$, Alice unveils her commitment to that index.  
Otherwise she announces failure and Bob rejects.\footnote{In the
ideal error-free case, if Alice knows $\eta$ precisely and both
parties honestly perform the protocol, failure is 
possible if and only if $q'>q$.} 
	
\item If Alice's unveiled commitment is indeed $x$, Bob accepts.   
Otherwise he rejects. 
	
\end{enumerate} 

We analyse the security of this protocol in the supplementary 
material.   We show the bound on $\epsilon_{\rm K}$ above holds for
any attacks by Bob, and that $\epsilon_{\rm C} \rightarrow 0$ 
for $q = \lfloor \frac{N}{d} \rfloor$ in the large $N$ limit. 
We also show that $\epsilon_{\rm S} \approx \frac{1}{d}$ if Alice is 
restricted to strategies in which she commits to $q$ classical
values chosen from $\{ 0 , \ldots , N+1 \}$ and unveils one of
these committed values.   
A full security analysis requires a complete analysis of general
quantum operations Alice could carry out to produce unveiling
data; we leave this for future work. 

\section{Conclusion} 
	
We have proven two no-go theorems demonstrating that even in the
relativistic setting there is no perfect KCEKQS protocol for quantum
states of finite dimension and bounding the evidence Alice can
supply. 
We have also described a 
new protocol involving quantum Bob-to-Alice communications
and relativistic signalling constraints, which appears to achieve a
significant improvement on existing protocols for large $d$. 
Although it is not zero knowledge for finite $d$, 
it reveals little extra information to Bob for large $d$.    
We conjecture that it is asymptotically CS-optimal, i.e. offers 
essentially optimal security against Alice.   
We anticipate that this protocol may be a valuable quantum 
cryptographic primitive in contexts where marginal revelations 
of information to Bob are acceptable.    	

{\bf Acknowledgments} \qquad This work was partially supported by an FQXi grant and by
Perimeter Institute for Theoretical Physics. Research at Perimeter
Institute is supported by the Government of Canada through Industry
Canada and by the Province of Ontario through the Ministry of Research
and Innovation.

\end{document}